# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

### v1

@@ Appendix D Proof of Theorem 3
Suppose we can find a set of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that $\min_{j\neq k}\|\mathcal{B}^{j}-\mathcal{B}^{k}\|_{F}\geq s$. According to [84, Theorem 4], when each element of $\bm{X}$ and $\mathcal{W}$ respectively follow $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, we have
@@
[…]
Next, we consider one construction for the sets of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that we can obtain a proper lower bound for $\min_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$ and a proper upper bound for $\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$. Given the equivalence between any two tensors $\mathcal{B}^{j_{1}}=[\bm{A}_{1}^{j_{1}},\mathcal{S},\bm{A}_{2}]$, $\mathcal{B}^{j_{2}}=[\bm{A}_{1}^{j_{2}},\mathcal{S},\bm{A}_{2}]$ and their canonical forms [48], the TT-SVD can be applied to transform these tensors into canonical formats $\mathcal{B}^{j_{1}}=[\widetilde{\bm{A}}_{1}^{j_{1}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ and $\mathcal{B}^{j_{2}}=[\widetilde{\bm{A}}_{1}^{j_{2}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ with $R(\widetilde{\mathcal{S}})R^{\rm H}(\widetilde{\mathcal{S}})={\bf I}_{N}$ and $\widetilde{\bm{A}}_{2}\widetilde{\bm{A}}_{2}^{\rm H}={\bf I}_{N}$. Given any $\delta>0$ and $N\geq C^{\prime}$ by [85, Lemma 5], we can construct a set of $\{\widetilde{\bm{A}}_{1}^{1},\dots,\widetilde{\bm{A}}_{1}^{n_{0}}\}$ with cardinality $n_{0}\geq\frac{1}{4}e^{\frac{LPN}{128}}$ such that: (1) $\|\widetilde{\bm{A}}_{1}^{j}\|_{F}=\delta$ holds for all $j=1,\dots,n_{0}$, (2) $\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta$ for all $j_{1},j_{2}\in[n_{0}],j_{1}\neq j_{2}$.
@@
Then we plug (D) into (50) and have
[…]
@@
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{UMN}{T}}\gamma\bigg{)}.$ (53)
@@
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{LPN+UMN}{2T}}\gamma\bigg{)}.$ (54)
Appendix E Proof of Theorem 5
@@
First, according to Theorem 4, we have
@@
$\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$
$=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N}^{D},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (57)
The last line follows that $\|\mathcal{H}\|_{F}^{2}=\|\bm{H}_{0}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A.
@@
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{0}\cdots n_{D})}\}$ with covering number
$n_{0}\cdots n_{D}\leq\bigg{(}\frac{4+\epsilon}{\epsilon}\bigg{)}^{\sum_{d=0}^{D}N_{d}N_{d+1}}$ (60)
for any $\mathcal{H}\in\mathbb{B}_{2N,\{\widetilde{\mathcal{S}}_{i}\}}^{D}$.
Denote by $A$ the value of (E), i.e.,
@@
[…]
$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$
Using $\mathcal{I}$ to denote the index set $[n_{0}]\times\cdots\times[n_{D}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{0},\dots,p_{D})\in\mathcal{I}$ such that
@@
[…]
@@
[…]
@@
[…]
@@
[…]
where the first equation follows the fact that $(l,k,u)$-th element of $[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}]$ can be decomposed into $\sum_{i=0}^{D}\bm{H}_{D}^{(p_{D})}(l,:)\widetilde{\mathcal{S}}_{D}(:,k,:)\bm{H}_{D-1}^{(p_{D-1})}\\ \cdots\bm{H}_{i+1}^{(p_{D_{i+1}})}\widetilde{\mathcal{S}}_{i+1}(:,k,:)(\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})})\widetilde{\mathcal{S}}_{i}(:,k,:)\widetilde{\bm{H}}_{i-1}\widetilde{\mathcal{S}}_{1}(:,k,:)\widetilde{\bm{H}}_{0}(:,u)$.
@@
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$ (66)
[…]

### v2

@@ Appendix F Proof of Theorem 3
Suppose we can find a set of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that $\min_{j\neq k}\|\mathcal{B}^{j}-\mathcal{B}^{k}\|_{F}\geq s$. According to [72, Theorem 4], when each element of $\bm{X}$ and $\mathcal{W}$ respectively follow $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, we have
@@
$\!\!\!\!\geq\!\!\!\!$ $\frac{s}{2}\bigg(1-\frac{\frac{T}{2\gamma^{2}}\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}^{2}+\log 2}{\log n}\bigg).$
Next, we consider one construction for the sets of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that we can obtain a proper lower bound for $\min_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$ and a proper upper bound for $\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$. Given the equivalence between any two tensors $\mathcal{B}^{j_{1}}=[\bm{A}_{1}^{j_{1}},\mathcal{S},\bm{A}_{2}]$, $\mathcal{B}^{j_{2}}=[\bm{A}_{1}^{j_{2}},\mathcal{S},\bm{A}_{2}]$ and their canonical forms [51], the TT-SVD can be applied to transform these tensors into canonical formats $\mathcal{B}^{j_{1}}=[\widetilde{\bm{A}}_{1}^{j_{1}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ and $\mathcal{B}^{j_{2}}=[\widetilde{\bm{A}}_{1}^{j_{2}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ with $R(\widetilde{\mathcal{S}})R^{\rm H}(\widetilde{\mathcal{S}})={\bf I}_{N}$ and $\widetilde{\bm{A}}_{2}\widetilde{\bm{A}}_{2}^{\rm H}={\bf I}_{N}$. Given any $\delta>0$ and $N\geq C^{\prime}$ by [99, Lemma 5], we can construct a set of $\{\widetilde{\bm{A}}_{1}^{1},\dots,\widetilde{\bm{A}}_{1}^{n_{0}}\}$ with cardinality $n_{0}\geq\frac{1}{4}e^{\frac{LPN}{128}}$ such that: (1) $\|\widetilde{\bm{A}}_{1}^{j}\|_{F}=\delta$ holds for all $j=1,\dots,n_{0}$, (2) $\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta$ for all $j_{1},j_{2}\in[n_{0}],j_{1}\neq j_{2}$.
@@
Then we plug (F) into (56) and have
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}$ $\!\!\!\!\geq\!\!\!\!$ $\frac{\delta}{2}\bigg(1-\frac{\frac{2T\delta^{2}}{\gamma^{2}}+\log 2}{c_{1}LPN}\bigg)$ (58)
$\!\!\!\!\geq\!\!\!\!$ $c_{2}\sqrt{\frac{LPN}{T}}\gamma,$
@@
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}=\Omega\bigg(\sqrt{\frac{UMN}{T}}\gamma\bigg).$ (59)
@@
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}=\Omega\bigg(\sqrt{\frac{LPN+UMN}{2T}}\gamma\bigg).$ (60)
Appendix G Proof of Theorem 5
@@
First, according to Theorem 4 in the main paper, we have
@@
$\hskip-14.22636pt\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$
$\hskip-14.22636pt=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N}^{D},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$\hskip-14.22636pt=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\hskip-31.2982pt\max_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}\hskip-36.98866pt|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (63)
The last line follows that $\|\mathcal{H}\|_{F}^{2}=\|\bm{H}_{0}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A of the main paper.
@@
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{0}\cdots n_{D})}\}$ with covering number $n_{0}\cdots n_{D}\leq(\frac{4+\epsilon}{\epsilon})^{\sum_{d=0}^{D}N_{d}N_{d+1}}$ for any $\mathcal{H}\in\mathbb{B}_{2N,\{\widetilde{\mathcal{S}}_{i}\}}^{D}$.
Denote by $A$ the value of (G), i.e.,
@@
$\!\!\!\!\hskip-14.22636pt=\!\!\!\!$ $\hskip-19.91684pt\operatorname*{\text{arg~max}}_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}\hskip-25.6073pt|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{1},\widetilde{\mathcal{S}}_{1},\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$
$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (67)
Using $\mathcal{I}$ to denote the index set $n_{0}]\times\cdots\times[n_{D}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{0},\dots,p_{D})\in\mathcal{I}$ such that
@@
$A$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$ (69)
@@
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
@@
$\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
@@
$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+\frac{A}{2},$
where the first equation follows the fact that $(l,k,u)$-th element of $\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}]$ can be decomposed into $\sum_{i=0}^{D}\bm{H}_{D}^{(p_{D})}(l,:)\widetilde{\mathcal{S}}_{D}(:,k,:)\bm{H}_{D-1}^{(p_{D-1})}\\ \cdots\bm{H}_{i+1}^{(p_{D_{i+1}})}\widetilde{\mathcal{S}}_{i+1}(:,k,:)(\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})})\widetilde{\mathcal{S}}_{i}(:,k,:)\widetilde{\bm{H}}_{i-1}\widetilde{\mathcal{S}}_{1}(:,k,:)\widetilde{\bm{H}}_{0}(:,u)$.
@@
$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=1}^{D-1}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg).$
Appendix H Error Analysis of channel estimation in the Multi-Hop IRS-Assisted MIMO Systems with special structures
In this section, we present the error analysis of channel estimation in multi-hop IRS-assisted MIMO systems under three special structures, analogous to the single-hop case.
$(i)$ First, when $\bm{B}_{d},d=0,\dots D$ are rank-$r_{d}$ matrices, by incorporating the covering number of low-rank matrices into Appendix G, we can derive
$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}(N_{d}+N_{d+1})r_{d})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
where $\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]\ \text{are unknown with }\text{rank}(\bm{B}_{d})=r_{d},d=0,\dots D,\ \text{and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$. Leveraging low-rank structures offers a potential avenue for improving error performance, as the recovery error may be reduced below eq. (14) in the main paper when the ranks $r_{d},d=0,\dots,D$ are small.
$(ii)$ Second, when the channel matrices exhibit sparsity in the angular domain, we assume the factorized form $\bm{B}_{d}=\bm{B}_{d1}\bm{\Sigma}_{d}\bm{B}_{d2},d=0,\dots D$, where $\bm{\Sigma}_{d}\in\mathbb{C}^{F_{d}\times F_{d}}$ is a diagonal matrix with $\|\bm{\Sigma}_{d}\|_{0}=s_{d}$, and $\bm{B}_{d1}\in\mathbb{C}^{N_{d+1}\times F_{d}}$, $\bm{B}_{d2}\in\mathbb{C}^{F_{d}\times N_{d}}$. Following a similar approach as in Appendix E, we obtain the error bound:
$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}s_{d}\log\big(\frac{F_{d}}{s_{d}}\big))\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
where $\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{d}=\bm{B}_{d1}\bm{\Sigma}_{d}\bm{B}_{d2},d=0,\dots D\ \text{where}\ \bm{\Sigma}_{d}\in\mathbb{C}^{F_{d}\times F_{d}}\ \text{is unknown}\text{with}\ \|\bm{\Sigma}_{d}\|_{0}=s_{d},\ \text{and}\ \bm{B}_{d1}\in\mathbb{C}^{N_{d+1}\times F_{d}},\bm{B}_{d2}\in\mathbb{C}^{F_{d}\times N_{d}}\ \text{are known, and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$. Note that the sparsity in the angular domain leads to a smaller recovery error compared to the non-sparse setting when $s_{d}\ll F_{d}$.
$(iii)$ Finally, when $\mathcal{S}_{i},i\in[D]$, are unknown, introducing the covering number of $\{\mathcal{S}_{i}\}_{i=1}^{D}$ into the proof in Appendix G, we can arrive at
$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
$\!\!\!\!=\!\!\!\!$ $O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1}+\sum_{d=1}^{D}KN_{d})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
where $\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{N_{D+1}\times K\times N_{0}}:\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d=0,\dots,D,\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are unknown}\}$. In this analysis, we assume that all IRSs are equipped with active elements. However, in practice, when adjustments are required for only a subset of IRSs, denoted as $\mathcal{S}_{i}$ for $i\in\Omega$, where $\Omega\subseteq\{1,\dots,D\}$, the recovery error can be further reduced. Specifically, the term $\sum_{d=1}^{D}KN_{d}$ is replaced by $\sum_{i\in\Omega}KN_{i}$.